OpenStudy (anonymous):

Lena bought a suit on sale for $555 . This price was 26% less than the original price. What was the original price? Help? please? I can't get an answer

jimthompson5910 (jim_thompson5910):

Let x = original price (Original Price) - (26% of original price) = Final Price (x) - (26% of x) = 555 x - 0.26x = 555 0.74x = 555 x = ???

OpenStudy (anonymous):

750?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

:D

jimthompson5910 (jim_thompson5910):

notice that 26% of 750 is 0.26*750 = 195 dollars

jimthompson5910 (jim_thompson5910):

this is the discount (and the amount you save) so the final price is 750 - 195 = 555, which is exactly what was given, so that checks the answer

OpenStudy (anonymous):

Thank you c:

jimthompson5910 (jim_thompson5910):

you're welcome